# FingerViP: Learning Real-World Dexterous Manipulation with Fingertip Visual Perception

### 23 Apr 2026 v1

@@ Abstract
The current practice of dexterous manipulation generally relies on a single wrist-mounted view, which is often occluded and limits performance on tasks requiring multi-view perception. In this work, we present FingerViP, a learning system that utilizes a visuomotor policy with fingertip visual perception for dexterous manipulation. Specifically, we design a vision-enhanced fingertip module with an embedded miniature camera and install the modules on each finger of a multi-fingered hand. The fingertip cameras substantially improve visual perception by providing comprehensive, multi-view feedback of both the hand and its surrounding environment. Building on the integrated fingertip modules, we develop a diffusion-based whole-body visuomotor policy conditioned on a third-view camera and multi-view fingertip vision, which effectively learns complex manipulation skills directly from human demonstrations. To improve view–proprioception alignment and contact awareness, each fingertip visual feature is augmented with its corresponding camera pose encoding and per-finger joint-current encoding. We validate the effectiveness of the multi-view fingetip vision and demonstrate the robustness and adaptability of FingerViP on various challenging real-world tasks, including pressing buttons inside a confined box, retrieving sticks from an unstable support, retrieving objects behind an occluding curtain, and performing long-horizon cabinet opening and object retrieval, achieving an overall success rate of $80.8\%$. All hardware designs and code will be fully open-sourced.
@@ I Introduction
Building on our fingertip design, we propose a learning-based framework to train a visuomotor policy directly from human demonstrations, enabling a systematic evaluation of the advantages of fingertip visual perception for dexterous manipulation. To address the challenges posed by dynamic, high-dimensional multi-fingertip inputs with continuously changing viewpoints, we introduce a transformer-based diffusion model [12] in which cross-attention captures semantic relationships between multi-fingertip image tokens and action tokens. Additionally, to preserve the spatial information and contact-related signals of fingertip vision, fingertip-view poses and per-finger joint currents are encoded and concatenated to visual features. By fusing fingertip (hand-centric) observations with third-view (global) context, the policy leverages complementary cues for both contact-rich interaction and scene-level awareness, facilitating robust policy learning from human demonstrations.
@@ III-B Vision-Enhanced Fingertip Module
Figure 3: Whole-Body Visuomotor Policy with Fingertip Visual Perception. (a) At time $t$, the policy observes five fingertip RGB images (gray) and one third-view image (pink), 20 hand joint currents, and 26 arm-hand joint angles. Fingertip camera poses are computed from hand forward kinematics. Visual features are extracted using a frozen vision encoder (CLIP), while camera poses, per-finger currents, and joint angles are embedded using MLP encoders and concatenated to construct the whole-body multi-modal perception tokens. (b) Conditioned on these tokens, a transformer-based diffusion model performs $K$ denoising steps to generate a horizon of $T$ whole-body action tokens, which are decoded by an MLP into 26-DoF joint commands.
@@ III-D Dexterous Visuomotor Policy
Policy Overview: Our policy is designed based on the Diffusion Policy [12], which infers joint-space actions via denoising diffusion conditioned on the latest two steps of observation data. Concretely, at each time step $t$, the policy takes both the previous observation $O_{t-1}$ and current observations $O_{t}$ as conditional inputs and predicts an $n$-step action sequence $A_{t}=\{a_{t},a_{t+1},\dots,a_{t+n}\}$, where only $m\leq n$ actions are executed. In our implementation, we set $n=16$ and $m=8$. The observation at time step $t$ is $O_{t}=\{I_{t},P_{t},C_{t},Q_{t}\}$, where $I_{t}\in\mathbb{R}^{6\times 224\times 224\times 3}$ consists of six RGB images from one third-view camera and five fingertip cameras, $P_{t}\in\mathbb{R}^{5\times(3+6)}$ are the camera poses, $C_{t}\in\mathbb{R}^{20\times 1}$ are whole-hand joint currents and $Q_{t}\in\mathbb{R}^{26\times 1}$ are joint angles of both arm and hand. The actions are taken from human teleoperation commands during training and each $a_{t}$ is represented as 26 joint angles.
Coordinating Fingertip Views With Positional and Joint Current Encoding. The robot’s motion continuously changes each fingertip camera’s pose and, consequently, its observing region, yielding a much larger observation space than static views (e.g., a fixed top-down camera). To better coordinate these moving views and improve policy learning efficiency, we introduce view-dependent positional encoding for each fingertip camera. Specifically, we use Pinocchio to parse the hand URDF, identify individual links of the robot hand, and compute each fingertip camera pose in the hand base frame via the forward kinematics model $\mathcal{F}$ from the current hand joint angles $Q^{h}_{t}\subset Q_{t}$. Let $^{B}\mathbf{T}_{t}\in SE(3)$ denote the transformation from the hand base frame $B$ to camera frame $c$ at time step $t$:
@@
Transformer-based Multi-view Semantic Correspondence Learning. To efficiently extract task-relevant cues from five dynamically moving fingertip cameras and a fixed third-view camera, a pretrained vision encoder and a transformer decoder are used to extract and aggregate semantics across multiple views. Prior works [68, 13, 66] have shown that pretrained vision foundation models such as CLIP [54] and DINO [7] provide strong semantic representations and improve performance in visually complex robotic manipulation tasks. Therefore, a CLIP-pretrained model is deployed as the image encoder in this policy. Specifically, the raw $640\times 480$ images are center-cropped to $224\times 224$ and processed by a frozen CLIP ViT-B/16 encoder [15], yielding a $768$-D class token feature per view. Features from the five fingertip cameras are linearly projected to $384$-D using a shared linear projector, while the third-view feature is projected with a separate (view-specific) linear layer. Each projected $384$-D visual feature is then concatenated with its corresponding camera-pose embedding and finger-current embedding to form a $768$-D token per camera. Robot joint angles are embedded to $768$-D and appended, producing a $(6+26)\times 768$ token sequence at time $t$. For action generation, we adopt a time-series diffusion transformer architecture [12] to iteratively denoise noisy action samples $A_{t}^{k}$, where $k$ denotes the denoising step. At denoising step $k$, noisy action samples $A_{t}^{k}$ are fed as input tokens, while the observation tokens condition the model through multi-head cross-attention [61]. This design efficiently fuses multi-view vision and proprioception and learns the correspondence between observations and joint-space actions.

### 05 May 2026 v2

@@ Abstract
The current practice of dexterous manipulation generally relies on a single wrist-mounted view, which is often occluded and limits performance on tasks requiring multi-view perception. In this work, we present FingerViP, a learning system that utilizes a visuomotor policy with fingertip visual perception for dexterous manipulation. Specifically, we design a vision-enhanced fingertip module with an embedded miniature camera and install the modules on each finger of a multi-fingered hand. The fingertip cameras substantially improve visual perception by providing comprehensive, multi-view feedback of both the hand and its surrounding environment. Building on the integrated fingertip modules, we develop a diffusion-based whole-body visuomotor policy conditioned on a third-view camera and multi-view fingertip vision, which effectively learns complex manipulation skills directly from human demonstrations. To improve view-proprioception alignment and contact awareness, each fingertip visual feature is augmented with its corresponding camera pose encoding and per-finger joint-current encoding. We validate the effectiveness of the multi-view fingertip vision and demonstrate the robustness and adaptability of FingerViP on various challenging real-world tasks, including pressing buttons inside a confined box, retrieving sticks from an unstable support, retrieving objects behind an occluding curtain, and performing long-horizon cabinet opening and object retrieval, achieving an overall success rate of $80.8\%$. All hardware designs and code will be fully open-sourced.
@@ I Introduction
Building on our fingertip design, we propose a learning-based framework to train a visuomotor policy directly from human demonstrations, enabling a systematic evaluation of the advantages of fingertip visual perception for dexterous manipulation. To address the challenges posed by dynamic, high-dimensional multi-fingertip inputs with continuously changing viewpoints, we use a transformer-based diffusion model [11] that conditions action denoising on fingertip visual tokens and robot proprioception. Additionally, to preserve the spatial information and contact-related signals of fingertip vision, fingertip-view poses and per-finger joint currents are encoded and concatenated to visual features. By fusing fingertip (hand-centric) observations with third-view (global) context, the policy leverages complementary cues for both contact-rich interaction and scene-level awareness, facilitating robust policy learning from human demonstrations.
@@ III-B Vision-Enhanced Fingertip Module
Figure 3: FingerViP Whole-Body Policy with Fingertip Visual Perception. (a) FingerViP collects five fingertip RGB images (gray) and one third-view image (pink) which provides global scene context, 20 hand joint currents, and 26 arm–hand joint angles at each time step. (b) Finger joint currents and fingertip camera poses derived from the hand kinematic model are encoded to provide contact-related cues and capture fingertip-view geometric information. Visual observations from the five fingertip cameras and one third-view camera are encoded using a vision foundation model. The resulting visual, geometric, current, and proprioceptive features are then concatenated to construct whole-body multi-modal perception tokens. (c) A transformer-based diffusion model (inspired by [11, 67, 61]) uses this token sequence to predict a future $26$-DoF joint-command trajectory for the robot arm and dexterous hand.
@@ III-D Dexterous Visuomotor Policy
Policy Overview: We formulate whole-body action prediction as a conditional denoising process following Diffusion Policy [11, 67]. At each time step $t$, the policy uses a short observation history, consisting of $O_{t-1}$ and $O_{t}$, to generate a future joint-command trajectory $A_{t}=\{a_{t},a_{t+1},\ldots,a_{t+n-1}\}$. Following a receding-horizon control strategy, only the first $m$ actions are applied to the robot before the policy is queried again, where $m\leq n$. Following common practice, we set $m=8$ and $n=16$ in our experiments. The observation $O_{t}$ contains both visual and proprioceptive information: $O_{t}=\{I_{t},P_{t},C_{t},Q_{t}\}$. Here, $I_{t}$ denotes the RGB observations from one third-view camera and five fingertip-mounted cameras, with the resolution of $244\times 224$. The hand-centric poses $P_{t}$ of each fingertip camera, and proprioception include the whole-hand joint-current vector $C_{t}\in\mathbb{R}^{20\times 1}$, and the $26$ arm-hand joint positions $Q_{t}$ are provided. The action $a_{t}$ is represented as a 26-dimensional joint-space command covering the 6-DoF robot arm and the 20-DoF dexterous hand. During training, these action targets are obtained from teleoperation demonstrations.
Coordinating Fingertip Views With Positional and Joint Current Encoding. In FingerViP, the fingertip cameras are rigidly mounted on articulated fingers, so their image observations are naturally tied to the current hand configuration. The visual feature extracted from a fingertip image is therefore not only view-specific but also finger-state-dependent. To expose both the view-dependent image context and mechanical context to the policy, we attach a camera-pose embedding to each fingertip visual token, then aggregate it with finger joint currents. To be more specific, inspired by the positional encoding strategy in [67], we compute the pose of each fingertip camera with respect to the hand base using forward kinematics and linearly project the camera pose to an embedding for each fingertip view. Specifically, we use Pinocchio to parse the hand URDF, identify individual links of the robot hand, and compute each fingertip camera pose in the hand base frame via the forward kinematics model $\mathcal{F}$ from the current hand joint angles $Q^{h}_{t}\subset Q_{t}$. Let $^{B}\mathbf{T}_{t}\in SE(3)$ denote the transformation from the hand base frame $B$ to camera frame $c$ at time step $t$:
@@
Transformer-based Diffusion Policy Learning. For action generation, we use a transformer-based diffusion policy [11] to predict a future joint-space action trajectory. During each denoising iteration $k$, the decoder refines a noisy action trajectory $A_{t}^{k}$ by querying the observation-token sequence through cross-attention [60]. The resulting denoised trajectory is used as the predicted future joint-command sequence.
The observation tokens are constructed by combining multi-view visual features and proprioceptive embeddings. Specifically, we use a frozen CLIP-pretrained ViT-B/16 encoder [53, 14] to extract visual features from the five fingertip cameras and the third-view camera. Prior robotic manipulation works have shown that vision foundation models such as CLIP can provide effective representations for challenging vision-based manipulation tasks [68, 12, 65, 67]. The five fingertip-view features are projected by a shared linear layer, whereas the third-view feature is processed by a separate (view-specific) projection layer. For each fingertip camera, the projected visual feature is concatenated with the corresponding camera-pose embedding and finger-current embedding to form a fingertip observation token. The third-view feature is used as a global visual token. Finally, the embedded robot joint positions are appended to the visual tokens, producing a $(6+26)\times 768$ observation-token sequence at time step $t$ for action prediction.